please help me solve this problem:
in thirty years, Roselle will be 8 times as old as she was 12 years ago. how old is she now?

Answer by boilpoil(127)

Let x be her age now

x+30=8(x-12)
x+30=8x-96
-7x=-126
x=18(years old)